# Flowr.root – A flow matching based foundation model for joint multi-purpose structure-aware 3D ligand generation and affinity prediction

### v2

@@ 4.4 Quantum Mechanical Calculations on Protein-Ligand Complexes
To benchmark Flowr.root’s affinity prediction head, we perform quantum mechanical (QM) calculations on protein-ligand complexes and correlate the results with predicted affinities. Calculations were conducted using ULYSSES (Menezes and Popowicz, 2022), a general-purpose semi-empirical library offering multiple molecular Hamiltonians. We employ GFN2-xTB (Bannwarth et al., 2019) with ALPB aqueous solvation (Ehlert et al., 2021). GFN2-xTB is a simplified density functional method that parametrizes interparticle interaction integrals against reference data, significantly accelerating calculations (Bannwarth et al., 2019)—a critical advantage for large atomic systems such as protein-ligand complexes. Additionally, GFN2-xTB provides a balanced treatment of electrostatic interactions and incorporates state-of-the-art dispersion corrections, enabling accurate descriptions of non-bonded complexes.
Benchmark calculations on protein complexes extracted from the Schrödinger FEP dataset already contained protonated protein structures. For these cases, Flowr.root-generated ligands had to be protonated using open babel (O’Boyle et al., 2011). Calculations were run on the full complexes, water molecules were excluded and solvation was treated only implicitly, except when otherwise stated. In the case of the kinase selectivity test, structures were extracted directly from the PDB and processed with open babel. This includes protonation of the protein pockets and of all ligands, along with their predicted total charges.
@@ 5 Results
Next, to benchmark Flowr.root for pocket-conditional 3D molecule generation, we use the widely adopted CrossDocked2020 dataset Francoeur et al. (2020), despite its prevalences (Cremer et al., 2025), and compare against recent state-of-the-art diffusion- and flow-based generative models, namely Pocket2Mol, DiffSBDD, TargetDiff, DrugFlow, PILOT and Flowr. For a more challenging and meaningful evaluation of generative SBDD models, we further evaluate Flowr.root on the recently proposed Spindr dataset and compare against Flowr and the diffusion-based Pilot model Cremer et al. (2024).
@@
We further demonstrate Flowr.root’s practical utility through domain adaptation via fine-tuning on project-specific in-house data and validate the model’s joint structure-affinity predictions through case studies on kinase selectivity (CK2/CLK3), ligand optimization (TYK2, ER), and quantum mechanical validation.
@@ 5.1 Unconditional Molecule Generation: Geom-Drugs
Flowr.root achieves the second highest RDKit-validity of and a PoseBusters-validity of , surpassing all other models significantly including FlowMol3 ( vs. ). These results indicate that Flowr.root generates molecules that are both chemically valid and geometrically plausible. It is worth noting that RDKit validity only measures whether or not a molecule can be processed by RDKit, while PoseBusters-validity is a comprehensive measure of 3D plausibility (Buttenschoen et al., 2024).
@@ 5.5 Affinity prediction: Schrodinger FEP+ and OpenFE
We further validate Flowr.root on the OpenFE IndustryBenchmark dataset, which is a curated subset of the Schrodinger FEP+ dataset. The top left panel of Figure 6 presents the scatter plot for Flowr.root-predicted pKd values versus experimental binding affinities. As with the Schrodinger FEP+ benchmark, the plot clearly shows a strong correlation with minimal deviation, suggesting that Flowr.root can predict binding affinities across a broad range of protein-ligand complexes. In the top right panel we show the performance of Flowr.root using various error and correlation metrics in comparison to OpenFE, FEP+, AEV-PLIG and Boltz-2. In the bottom left panel, we show the combined affinity prediction (median over pKd, pKi, and pIC.
@@ 5.7 Case studies
Second, we benchmark the framework’s conditional ligand generation performance against quantum mechanical binding energy calculations using TYK2 kinase and ER as model systems. These validation studies assess the correlation between Flowr.root-predicted binding affinities and computationally demanding QM binding energies, while providing mechanistic insights into the structural features that govern binding affinity in the generated ligands.
@@ 5.8 Ligand Selectivity: CK2 and CLK3
The human kinome consists of over 500 kinases (Manning et al., 2002) that share conserved catalytic domain structures, particularly in their ATP-binding sites, where most competitive inhibitors bind. This structural homology, while functionally important for cellular signaling pathways, presents a major challenge for kinase selectivity—the ability of small molecule inhibitors to achieve potent inhibition of specific kinase targets while avoiding off-target interactions with structurally similar kinases. To demonstrate the capability of our framework for multi-objective optimization, we perform a selectivity study targeting selectivity between CK2 (on-target, PDB: 3pe1) and CLK3 (off-target, PDB: 6khf).
We compare two optimization strategies using importance sampling: joint optimization that simultaneously maximizes the predicted binding affinity on CK2 while minimizing CLK3 affinity formulated as dual-objective, versus the single optimization which only maximizes CK2. We generated ligands for each set using the protein pockets from PDB 3pe1 (Battistutta et al., 2011) and 6khf (Lee et al., 2019) after rotational alignment.
@@
The second mechanism is related to interactions with H160. In previous work (Grygier et al., 2023), we identified H160 as a critical residue conferring sub-nM affinity of the inhibitor silmitasertib to CK2. Here too, we observe that Flowr.root explores interactions with this residue to achieve selectivity. On the CK2-selective binder, Flowr.root orients the ligand’s phenyl substituent to optimize a T-stack interaction. On the CLK3-selective binder, a similar substituent is rotated to yield a short, repulsive contact. This is possible because the side chain of E287, the residue in a similar position in CLK3, points in another direction.
@@
5.9 TYK2 and ER
As shown in previous sections, Flowr.root generates novel chemical matter conditioned on a protein pocket. The generated molecules are chemically reasonable and are not heavily strained, unlike previous generative models. Additionally, the predicted affinity data reliably reproduces experimental distributions. In this section, we benchmark how the whole Flowr.root workflow behaves on the task of conditional ligand generation. To this end, we chose the TYK2 kinase and ER as targets from the Schrodinger FEP+ dataset.
@@ 5.9 TYK2 and ER
We first look at the example of the TYK2 kinase (Fig. 10, upper block). We find a good correlation between the QM-calculated binding energies and the Flowr.root-predicted median affinities, with an R2 of , a Pearson correlation coefficient of , a Kendall of , RMSE of , and RAE of . From a chemical viewpoint, it is more interesting to compare the structures of the best and worst Flowr.root-predicted ligands. We see that, despite some common features, the model generalizes effectively the hinge-binding motif, and the solvent-exposed residue of the ligand. Importantly, all ligands retained the hydrogen bond acceptor to V981, a critical hinge residue already in the original ligand structure. Binding to a kinase’s hinge is a very important in kinase inhibition, kept by Flowr.root-generated chemical matter. Interestingly, no ligand improved the hydrogen-bond geometry with P982, despite some variations in the chemotypes exhibited by the hinge-binding heads. The major differences take place in how specific interactions are exploited. Particularly critical for the best binder in this series is the -stack interaction established with Y980. But also the fact that this ligand protects the least lipophilic halogens from solvent exposure. We also calculated ligand deformation energies, which for this series lie between kcal/mol and kcal/mol.
@@ 6 Conclusion
In this work, we introduced Flowr.root, a foundational framework for structure-based generative chemistry that unifies multi-mode ligand generation with binding affinity prediction and domain adaptation. Built on an -equivariant flow matching backbone, Flowr.root supports de novo design, interaction-guided generation, and fragment-based elaboration within a single architecture to allow applications from hit identification to lead optimization. By coupling structural prediction with dedicated affinity and confidence heads, the model directly integrates potency estimation into the generative trajectory. Our three-stage training paradigm—large-scale pretraining on billions of ligands and millions of protein–ligand complexes, refinement on curated higher-fidelity datasets, and project-specific adaptation—enables the model to capture broad chemical diversity while remaining efficiently adaptable to specialized structure-activity landscapes. This design philosophy reflects a fundamental insight: rather than expecting universal generalization, effective computational tools must function as continuously evolving companions that refine their understanding through sustained interaction with project-specific data.
@@
At the same time, several limitations remain. The reliance on public datasets, despite extensive curation, introduces biases from noisy affinity measurements, overrepresented protein families, and uneven chemical diversity. While higher-fidelity datasets partially mitigate these issues, their limited scale constrains model calibration, particularly for underexplored targets. Moreover, project-specific adaptation, while powerful, requires carefully chosen objectives and sufficient assay data; otherwise, models risk overfitting to narrow distributions or encoding spurious preferences. Also, Flowr.root requires the binding pocket to be known and preferably in a holo state. We leave the flexibility of protein pockets to future work. Finally, although Flowr.root incorporates an affinity head, its predictions remain approximations and require experimental validation.
@@
Looking forward, several directions appear promising. Expanding project-level adaptation to include reinforcement learning or active learning with experimental feedback may enable continuous refinement during discovery campaigns. Enhancing synthetic accessibility estimation and coupling to reaction-based generative models would improve downstream feasibility. Finally, bridging generative chemistry with physics-based scoring and simulation may yield hybrid frameworks that combine the scalability of learning-based models with the rigor of molecular physics.
Overall, Flowr.root provides a step towards foundational, adaptable generative models for structure-based drug discovery. By explicitly unifying multi-mode ligand generation, affinity prediction, and project-specific domain adaptation, it lays the groundwork for future systems that can generalize broadly, adapt flexibly, and integrate well into medicinal chemistry workflows.
@@ A.1 KIBA-3D
The KIBA dataset was originally introduced by Tang et al. [Tang et al., 2014b] as a large-scale benchmark for kinase–ligand binding affinity prediction, integrating heterogeneous bioactivity readouts into a standardized KIBA score. The original dataset contained chemical compounds, kinase targets and over interaction records. While comprehensive, the original KIBA dataset was highly scattered, due to an imbalance between the number of ligand-kinase pairs and associated measurements. To improve the density and ensure more reliable evaluation, subsequent studies discarded the ligands and targets with less than 10 interaction records. This procedure reduced the dataset to ligands and kinase targets. For this work, we prepared the 3D extension of KIBA (KIBA-3D) based on kinase targets with experimentally solved structures deposited in the PDB. Figure shows examples of our docking results, illustrating that Glide could reproduce near-native binding poses for selected kinase inhibitors with conserved bidentate hydrogen bond pattern in the hinge region (e.g., PDBID:2XB7-CHEMBL461876, glide gscore , PDBID:6LVM-CHEMBL1968590, glide gscore ). The KIBA-3D dataset serves as a large-scale, structure-consistent benchmark for evaluating generative and predictive models.
@@ A.1.2 Docking
We performed docking with the Maestro suite (Schrödinger Release 2023-2: Maestro, Schrödinger, LLC, New York, NY, 2023) to generate consistent protein–ligand complexes across the kinase targets. Small molecules were first optimized with LigPrep. All the ligands were processed. The dataset covers a broad chemical space, with molecular weights ranging from to Da (median Da). Protein structures were processed using the Protein Preparation Workflow, including protonation with Epik and propka [Johnston et al., 2023] and restrained minimization with the OPLS4 force field [Sastry et al., 2013]. Missing residues were built with Prime. Docking grids were centered on the kinase domain’s hinge region, covering the canonical ATP-binding pocket used by first-type kinase inhibitors. Docking was performed with the Glide SP protocol [Yang et al., 2021], with no constraints applied, generating one pose per ligand.

### v3

@@ 4.4 Quantum Mechanical Calculations on Protein-Ligand Complexes
To benchmark Flowr.root’s affinity prediction head, we perform quantum mechanical (QM) calculations on protein-ligand complexes and correlate the results with predicted affinities. Calculations were conducted using ULYSSES (Menezes:ULYSSES), a general-purpose semi-empirical library offering multiple molecular Hamiltonians. We employ GFN2-xTB (xtb_gfn2) with ALPB aqueous solvation (alpb). GFN2-xTB is a simplified density functional method that parametrizes interparticle interaction integrals against reference data, significantly accelerating calculations (xtb_gfn2)—a critical advantage for large atomic systems such as protein-ligand complexes. Additionally, GFN2-xTB provides a balanced treatment of electrostatic interactions and incorporates state-of-the-art dispersion corrections, enabling accurate descriptions of non-bonded complexes.
Benchmark calculations on protein complexes extracted from the Schrödinger FEP dataset already contained protonated protein structures. For these cases, Flowr.root-generated ligands had to be protonated using open babel (obabelpaper). Calculations were run on the full complexes, water molecules were excluded and solvation was treated only implicitly, except when otherwise stated. In the case of the kinase selectivity test, structures were extracted directly from the PDB and processed with open babel. This includes protonation of the protein pockets and of all ligands, along with their predicted total charges.
@@ 5 Results
Next, to benchmark Flowr.root for pocket-conditional 3D molecule generation, we use the widely adopted CrossDocked2020 dataset dataset:crossdocked, despite its prevalences (model:flowr), and compare against recent state-of-the-art diffusion- and flow-based generative models, namely Pocket2Mol, DiffSBDD, TargetDiff, DrugFlow, PILOT and Flowr. For a more challenging and meaningful evaluation of generative SBDD models, we further evaluate Flowr.root on the recently proposed Spindr dataset and compare against Flowr and the diffusion-based Pilot model model:pilot.
@@
We further demonstrate Flowr.root’s practical utility through domain adaptation via fine-tuning on project-specific in-house data and validate the model’s joint structure-affinity predictions through case studies on kinase selectivity (CK2/CLK3), ligand optimization (TYK2, ER, BACE1), and quantum mechanical validation.
@@ 5.1 Unconditional Molecule Generation: Geom-Drugs
Flowr.root achieves the second highest RDKit-validity of and a PoseBusters-validity of , surpassing all other models significantly including FlowMol3 ( vs. ). These results indicate that Flowr.root generates molecules that are both chemically valid and geometrically plausible. It is worth noting that RDKit validity only measures whether or not a molecule can be processed by RDKit, while PoseBusters-validity is a comprehensive measure of 3D plausibility (benchmarking:posebusters).
@@ 5.5 Affinity prediction: Schrodinger FEP+ and OpenFE
We further validate Flowr.root on the OpenFE IndustryBenchmark dataset, which is a curated subset of the Schrodinger FEP+ dataset 111Data from (08/2025) https://github.com/OpenFreeEnergy/IndustryBenchmarks2024222Results from (08/2025) https://blog.omsf.io/the-free-energy-of-everything-benchmarking-openfe. The top left panel of Figure 6 presents the scatter plot for Flowr.root-predicted pKd values versus experimental binding affinities. As with the Schrodinger FEP+ benchmark, the plot clearly shows a strong correlation with minimal deviation, suggesting that Flowr.root can predict binding affinities across a broad range of protein-ligand complexes. In the top right panel we show the performance of Flowr.root using various error and correlation metrics in comparison to OpenFE, FEP+, AEV-PLIG and Boltz-2. In the bottom left panel, we show the combined affinity prediction (median over pKd, pKi, and pIC.
@@ 5.7 Case studies
Second, we benchmark the framework’s conditional ligand generation performance against quantum mechanical binding energy calculations using TYK2 kinase, ER and BACE1 as model systems. These validation studies assess the correlation between Flowr.root-predicted binding affinities and computationally demanding QM binding energies, while providing mechanistic insights into the structural features that govern binding affinity in the generated ligands.
@@ 5.8 Ligand Selectivity: CK2 and CLK3
The human kinome consists of over 500 kinases (doi:10.1126/science.1075762) that share conserved catalytic domain structures, particularly in their ATP-binding sites, where most competitive inhibitors bind. This structural homology, while functionally important for cellular signaling pathways, presents a major challenge for kinase selectivity—the ability of small molecule inhibitors to achieve potent inhibition of specific kinase targets while avoiding off-target interactions with structurally similar kinases. To demonstrate the capability of our framework for multi-objective optimization, we perform a selectivity study targeting selectivity between CK2 (on-target, PDB: 3pe1) and CLK3 (off-target, PDB: 6khf).
We compare two optimization strategies using importance sampling: joint optimization that simultaneously maximizes the predicted binding affinity on CK2 while minimizing CLK3 affinity formulated as dual-objective, versus the single optimization which only maximizes CK2. We generated ligands for each set using the protein pockets from PDB 3pe1 (Battistutta2011-dh) and 6khf (Lee2019-fd) after rotational alignment.
@@
The second mechanism is related to interactions with H160. In previous work (menezes:silmitasertib), we identified H160 as a critical residue conferring sub-nM affinity of the inhibitor silmitasertib to CK2. Here too, we observe that Flowr.root explores interactions with this residue to achieve selectivity. On the CK2-selective binder, Flowr.root orients the ligand’s phenyl substituent to optimize a T-stack interaction. On the CLK3-selective binder, a similar substituent is rotated to yield a short, repulsive contact. This is possible because the side chain of E287, the residue in a similar position in CLK3, points in another direction.
@@
5.9 TYK2, ER, and BACE1
As shown in previous sections, Flowr.root generates novel chemical matter conditioned on a protein pocket. The generated molecules are chemically reasonable and are not heavily strained, unlike previous generative models. Additionally, the predicted affinity data reliably reproduces experimental distributions. In this section, we benchmark how the whole Flowr.root workflow behaves on the task of conditional ligand generation. To this end, we chose the TYK2 kinase, ER, and BACE1 as targets from the Schrodinger FEP+ dataset.
@@ 5.9 TYK2, ER, and BACE1
We first look at the example of the TYK2 kinase (Fig. 10, upper block). We find a good correlation between the QM-calculated binding energies and the Flowr.root-predicted median affinities, with an R2 of , a Pearson correlation coefficient of , a Kendall of , RMSE of , and RAE of . From a chemical viewpoint, it is more interesting to compare the structures of the best and worst Flowr.root-predicted ligands. We see that, despite some common features, the model generalizes effectively the hinge-binding motif, and the solvent-exposed residue of the ligand. Importantly, all ligands retained the hydrogen bond acceptor to V981, a critical hinge residue already in the original ligand structure. Binding to a kinase’s hinge is very important in kinase inhibition, kept by Flowr.root-generated chemical matter. Interestingly, no ligand improved the hydrogen-bond geometry with P982, despite some variations in the chemotypes exhibited by the hinge-binding heads. The major differences take place in how specific interactions are exploited. Particularly critical for the best binder in this series is the -stack interaction established with Y980. But also the fact that this ligand protects the least lipophilic halogens from solvent exposure. We also calculated ligand deformation energies, which for this series lie between kcal/mol and kcal/mol.
@@
Finally, we also analysed a benchmark on BACE1, where we obtained the best correlation of this series between Flowr.root-predicted affinities and the QM calculated ones. This is seen in the lowest RAE and RMSE coupled with the highest R2, Pearson correlation coefficient, and Kendall of this series of benchmarks. We see that in all cases the ligands retain a water-mediated hydrogen bond with F108, which results in high affinities overall. The major differences are in the nature of the nitrogen-containing group near Y71, which seems to favour more lipophilic groups, but also the substituents in the aromatic ring T-stacking the residue W115. We note that the T-stack interaction is favoured by electron-rich phenyl rings, which is not the case of the worst binder in the series. This seems then to be one of the major effects controlling affinity to BACE1.
@@ 6 Conclusion
In this work, we introduced Flowr.root, a foundational framework for structure-based generative chemistry that unifies multi-mode ligand generation with binding affinity prediction and domain adaptation within a single E(3)-equivariant flow matching architecture. The framework supports de novo design, interaction-guided generation, and fragment-based elaboration, enabling applications spanning from early-stage hit identification to late-stage lead optimization. By coupling structural prediction with dedicated affinity and confidence heads, Flowr.root directly integrates potency estimation into the generative trajectory, allowing property-guided molecule design through inference-time steering. Our three-stage training paradigm—large-scale pretraining on billions of ligands and millions of mixed-fidelity protein-ligand complexes, refinement on curated high-fidelity datasets, and project-specific adaptation—enables the model to capture broad chemical diversity while remaining efficiently adaptable to specialized structure-activity landscapes.
Our comprehensive evaluation demonstrates that Flowr.root advances the state of the art across multiple dimensions spanning from unconditional molecule to structure-aware ligand generation consistently outperforming recent diffusion- and flow-based approaches. For affinity prediction, Flowr.root achieves robust correlation with experimental measurements across diverse assay types, matching or exceeding the performance of computationally expensive physics-based and clearly outperforming ML-based methods while operating orders of magnitude faster. Critically, the model predictions maintain reliability across both public benchmarks and project-specific datasets when followed by targeted fine-tuning, demonstrating practical utility in real-world medicinal chemistry workflows.
Quantum mechanical validation studies on TYK2 kinase, ER, and BACE1 provide evidence of Flowr.root’s capacity to generate chemically meaningful ligands and binding poses. Strong correlations between model-predicted affinities and rigorous QM binding energies—coupled with mechanistic analysis revealing that generated ligands systematically exploit key structural motifs such as hinge-binding interactions, hydrogen bond geometries, and aromatic stacking—demonstrate that the model has internalized fundamental principles of protein-ligand interactions. Our kinase selectivity study between CK2 and CLK3 further illustrates that through affinity-guided generation, Flowr.root can help finding selectivity mechanisms. Overall, Flowr.root provides a concrete step toward the vision of establishing continuously evolving frameworks that refine their understanding through sustained interaction with project-specific data, offering a foundational yet adaptable architecture that generalizes broadly, and adapts flexibly to specialized domains integrating well into medicinal chemistry workflows.
@@
Nevertheless, important limitations remain. The reliance on public datasets, despite extensive curation, introduces biases from noisy affinity measurements, overrepresented protein families, and uneven chemical diversity. While higher-fidelity datasets partially mitigate these issues, their limited scale constrains model calibration, particularly for underexplored targets. Project-specific adaptation, though powerful, requires carefully chosen objectives and sufficient assay data; otherwise, models risk overfitting to narrow distributions. Additionally, Flowr.root requires the binding pocket to be known and preferably in a holo conformation, leaving the challenge of modeling protein pocket flexibility to future work. Finally, the predictions of Flowr.root remain approximations that cannot substitute for experimental validation.
@@
Looking forward, several directions appear promising. Expanding project-level adaptation to include reinforcement learning or active learning frameworks with both physics-based and experimental feedback may enable continuous refinement during discovery campaigns. Further, enhancing synthetic accessibility estimation and coupling to reaction-based generative models would improve downstream feasibility, ensuring that generated molecules are not only potent but also synthetically accessible.
Data availability
All datasets used in this study are publicly available and can be downloaded from Google Drive at https://drive.google.com/drive/folders/1NWpzTY-BG_9C4zXZndWlKwdu7UJNCYj8.
Code availability
The source code for Flowr.root is publicly available on GitHub at github.com/jule-c/flowr_root. The pretrained model checkpoint can be downloaded from Google Drive at https://drive.google.com/drive/folders/1NWpzTY-BG_9C4zXZndWlKwdu7UJNCYj8.
@@ A.1 KIBA-3D
The KIBA dataset was originally introduced by Tang et al. [tang2014] as a large-scale benchmark for kinase–ligand binding affinity prediction, integrating heterogeneous bioactivity readouts into a standardized KIBA score. The original dataset contained chemical compounds, kinase targets and over interaction records. While comprehensive, the original KIBA dataset was highly scattered, due to an imbalance between the number of ligand-kinase pairs and associated measurements. To improve the density and ensure more reliable evaluation, subsequent studies discarded the ligands and targets with less than 10 interaction records. This procedure reduced the dataset to ligands and kinase targets. For this work, we prepared the 3D extension of KIBA (KIBA-3D) based on kinase targets with experimentally solved structures deposited in the PDB. Figure shows examples of our docking results, illustrating that Glide could reproduce near-native binding poses for selected kinase inhibitors with conserved bidentate hydrogen bond pattern in the hinge region (e.g., PDBID:2XB7-CHEMBL461876, glide gscore , PDBID:6LVM-CHEMBL1968590, glide gscore ; see Fig. 16. The KIBA-3D dataset serves as a large-scale, structure-consistent benchmark for evaluating generative and predictive models.
@@ A.1.2 Docking
We performed docking with the Maestro suite (Schrödinger Release 2023-2: Maestro, Schrödinger, LLC, New York, NY, 2023) to generate consistent protein–ligand complexes across the kinase targets. Small molecules were first optimized with LigPrep. All the ligands were processed. The dataset covers a broad chemical space, with molecular weights ranging from to Da (median Da). Protein structures were processed using the Protein Preparation Workflow, including protonation with Epik and propka [johnston2023epik] and restrained minimization with the OPLS4 force field [sastry2013protein]. Missing residues were built with Prime. Docking grids were centered on the kinase domain’s hinge region, covering the canonical ATP-binding pocket used by first-type kinase inhibitors. Docking was performed with the Glide SP protocol [yang2021efficient], with no constraints applied, generating one pose per ligand.